Power Manufacturing has equipment that it purchased 7 years ago for $1,950,000. The equipment was used for a project that was intended to last for 9 years. However, due to low demand, the project is being shut down. The equipment was depreciated using the straight-line method and can be sold for $280,000 today. The company's tax rate is 40 percent. What is the aftertax salvage value of the equipment?

rev: 06_18_2018_QC_CS-129446

$310,667

$341,333

$218,667

$392,000

$280,000
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(1,950,000/9)=$216,666.67/year

Hence book value as on date of sale=Cost-Accumulated depreciation

=1,950,000-(216,666.67*7)

=$433,333.33

Hence loss on sale =$433,333.33-$280,000 =$153,333.33

Hence aftertax salvage value=Sale proceeds+(Tax rate*Loss on sale)

=280,000+(153,333.33*0.4)

=$341,333(Approx).